OpenStudy (anonymous):

3x - 47 = 7 How to find the slope ?

myininaya (myininaya):

the slope is undefined since it can be written in this form: x=a

OpenStudy (anonymous):

thanks !

myininaya (myininaya):

did you mean to have a y in there somewherE?

OpenStudy (anonymous):

The result is a vertical line at x = 18

OpenStudy (anonymous):

ugh . yes . it really is . 3x - 4y =7

myininaya (myininaya):

write in y=mx+b from and whatever your coeifficent of x is is your slope

myininaya (myininaya):

Solve for y: First you should subtract 3x on both sides giving you: -4y=-3x+7

myininaya (myininaya):

now divide both sides by -4 giving you: y= -3x/(-4)+7/(-4)=3x/4-7/4

myininaya (myininaya):

so since m=3/4 and m represents your slope then the slope is 3/4

OpenStudy (anonymous):

You can just divide the equation by the coefficient in front of the y. Then the coefficient in front of x will be the slope.

OpenStudy (anonymous):

thank you both so much (: